# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v5

@@ 2.1.2 Code-to-Code
Maniple (Parasaram et al., 2024), fill=cyan, fill opacity=0.2
[Code
Infilling, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
[, fill=cyan, fill opacity=0.2
StarCoder (Li et al., 2023h), Code LLaMA (Rozière et al., 2023), fill=cyan, fill opacity=0.2
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2]
[Text-to-Text, fill=Dandelion, fill opacity=0.4]
@@
where level=2text width=4.1em,font=,
where level=3text width=10em,font=,
where level=4text width=6em,font=,
where level=5text width=11em,font=,
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5
[Code-to-Code, fill=cyan, fill opacity=0.2
[Decompilation, fill=cyan, fill opacity=0.2
, fill=cyan, fill opacity=0.2
Neutron (Liang et al., 2021),
NeurDP (Cao et al., 2022)
, fill=cyan, fill opacity=0.2
BTC (Hosseini & Dolan-Gavitt, 2022),
SLaDe (Armengol-Estapé et al., 2023),
Nova+ (Jiang et al., 2023c),
LLM4Decompile (Tan et al., 2024)
, fill=cyan, fill opacity=0.2
[Deobfuscation, fill=cyan, fill opacity=0.2
DeGuard (Bichsel et al., 2016), Autonym (Vasilescu et al., 2017),
Debin (He et al., 2018), JSNeat (Tran et al., 2019), fill=cyan, fill opacity=0.2
DIRE (Lacomis et al., 2019), Artuso et al. (2021), fill=cyan, fill opacity=0.2
VarBERT (Banerjee et al., 2021), DIRECT (Nitin et al., 2021),
LmPa (Xu et al., 2023), fill=cyan, fill opacity=0.2
@@
DeepREL (Deng et al., 2022b), NNSmith (Liu et al., 2023e),
@@
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2]
[Text-to-Text, fill=Dandelion, fill opacity=0.4]
- Code search is a task similar to code retrieval, and differs from the later only in that the input is an existing code snippet, often in a different programming language from the target. Code search can be reformulated as finding clones of the query in the pool of targets, and is thus equivalent to clone detection to some extent.
- Code completion aims to complete a piece of code given its prefix, and remains to date one of the most popular applications of code language models in IDEs. This is essentially language modeling applied to code, and related technologies have been progressively introduced: n-gram, RNN, and Transformer. However, due to the structured nature of programming languages, many early works found grammar-aided statistical models to perform better (Bielik et al., 2016; Hellendoorn & Devanbu, 2017), and neural models only became dominant after 2018 (see Figure 4 for an intuitive overview.)
- Code translation aims to translate a piece of code (usually a function or method) into another programming language. The relation between code translation and cross-lingual code search is similar to the one between code synthesis and text-to-code retrieval, and SMT/MNT models have also been widely applied to this task. Code translation is an important technique in migrating old projects written in obsolete languages. However, we are yet to witness such applications, as the context window of even the most powerful language models are quite limited in the face of such projects. Malyala et al. (2023) provide a short survey on this task from the SE perspective.
- API mining refers to the process of finding similar APIs in different libraries, potentially in different programming languages. This task is closely related to API migration - a special case of code translation. API mining is traditionally tackled by computing similarity metrics between source and target APIs using information retrieval models, but as generative models become ever more capable, it is also worth exploring to directly generate the target API as a sequence-to-sequence task. Another closely related task is idiom mining (Allamanis & Sutton, 2014), where the objective is to discover commonly used code patterns, which exposes the potential need of new APIs (Sivaraman et al., 2022).
- Code repair, also known as bug fix, aims to fix a piece of buggy code. Like code translation, it is a traditional sequence-to-sequence generation task, and surveys are abundant on this topic (Gazzola et al., 2018; Monperrus, 2018; Zhong et al., 2022; Zhang et al., 2023d; Huang et al., 2023a). A generalization of code repair is code edit or code change, where the task is to model and predict code changes (such as Git commits), which are not necessarily about bug fixing.
- Code infilling is another recently proposed task, after fill-in-the-middle pretraining (Bavarian et al., 2022) became popular. It is a generalization of code completion, where not only the left context, but also the right context is given. However, it differs from cloze test in that the target of cloze test is only one token, while the target of code infilling can be an entire line or even multiple lines, which requires a decoder to generate autoregressive.
- Decompilation is the reverse process of compilation. In this task a model takes a compiled program - represented in assembly code or binary machine code - and aims to output the original source program in high-level languages such as C.
- Obfuscation refers to the process of renaming identifiers (e.g. variables, methods, and classes), for example to generic names like var_1, var_2 or x, y. It is an important technique in virus detection, intellectual property protection, and code size reduction (Collberg & Thomborson, 2002; Murad et al., 2010; Vasilescu et al., 2017). Deobfuscation refers to the reverse process, where meaningful identifier names are recovered from obfuscated programs. Obfuscation can be easily achieved statically, but deobfuscation has been a subject of more interest in recent years. It plays a significant role in decompiling or reverse engineering, and has also been adopted as a pretraining objective for code language models (Lachaux et al., 2021; Ding et al., 2022a; Liu et al., 2022d).
- Unit test generation aims to generate unit tests for a given program. Prior to the rise of Codex and other code LLMs, almost all works in this area employed non-neural methods (see Figure 5). In the age of LLMs, however, this task is ever more important, as research has shown that the current unit tests for evaluating LLMs’ program synthesis capability may be insufficient (Liu et al., 2023f). Wang et al. (2023a) provide a comprehensive survey on software testing with LLMs.
{forest}
for tree=
forked edges,
draw,
rounded corners,
grow=east,
anchor=base west,
anchor=center,
reversed=true,
l sep=0.2cm,
where level=0font=,
where level=1text width=4.9em,font=,
where level=2text width=3.5em,font=,
where level=3text width=7em,font=,
where level=4text width=9em,font=,
where level=5text width=11em,font=,
for tree=
if level=0l sep+=0.15cm,
if level=1l sep+=0.15cm,
if level=2l sep+=0.03cm,
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5]
[Code-to-Code, fill=cyan, fill opacity=0.4]
[Code-to-Text, fill=Orange, fill opacity=0.2
[Code Summarization, fill=Orange, fill opacity=0.2
CloCom (Wong et al., 2015), fill=Orange, fill opacity=0.2
CODE-NN (Iyer et al., 2016), DeepCom (Hu et al., 2018a),
TL-CodeSum (Hu et al., 2018b), Code2Seq (Alon et al., 2019a),
Fernandes et al. (2019), Wan et al. (2018),
AST-AttendGRU (LeClair et al., 2019), Wei et al. (2019),
Haque et al. (2020), DMACOS (Xie et al., 2021),
Bansal et al. (2021),
CoCoSUM (Wang et al., 2021d),
MLCS (Zhou et al., 2023c), fill=Orange, fill opacity=0.2
TranS3 (Wang et al., 2020e), Ahmad et al. (2020),
Corder (Bui et al., 2021b), SiT (Wu et al., 2021),
SG-Trans (Gao et al., 2023c),
Codex-D (Chen et al., 2021b),
M2TS (Gao & Lyu, 2022), AST-Trans (Tang et al., 2022),
GTrans (Kuang et al., 2022), CoSS (Shi et al., 2023a),
Sun et al. (2023b), Yuan et al. (2023a),
use-seq (Su & McMillan, 2023a),
Su & McMillan (2023b),
CSA-Trans (Oh & Yoo, 2024),
Haldar & Hockenmaier (2024)
, fill=Orange, fill opacity=0.2
[Code
Review, fill=Orange, fill opacity=0.2
Review Bot (Balachandran, 2013), Tricorder (Sadowski et al., 2015),
ClusterChanges (Barnett et al., 2015), CommentFinder (Hong et al., 2022), fill=Orange, fill opacity=0.2
DCR (Gupta & Sundaresan, 2018), DeepReview (Li et al., 2019a),
DACE (Shi et al., 2019), CORE (Siow et al., 2020), fill=Orange, fill opacity=0.2
Tufano et al. (2021b), Hellendoorn et al. (2021),
Tufano et al. (2022b), CodeReviewer (Li et al., 2022j),
AUGER (Li et al., 2022b), crBERT (Yin et al., 2023),
LLaMA-Reviewer (Lu et al., 2023a), fill=Orange, fill opacity=0.2
[Identifier Prediction, fill=Orange, fill opacity=0.2
JSNice (Raychev et al., 2015), Pigeon (Alon et al., 2018),
HeMa (Jiang et al., 2019), fill=Orange, fill opacity=0.2
Allamanis et al. (2015), Allamanis et al. (2016b),
GGNN (Allamanis et al., 2018), Code2Vec (Alon et al., 2019b),
Code2Seq (Alon et al., 2019a), Fernandes et al. (2019),
HIER (Xu et al., 2019b), Liu et al. (2019a),
MNire (Nguyen et al., 2020), DeepName (Li et al., 2021b),
DMACOS (Xie et al., 2021), NamPat (Yang et al., 2022), fill=Orange, fill opacity=0.2
GTNM (Liu et al., 2022b), GTrans (Kuang et al., 2022),
Mario (Wang et al., 2023c), fill=Orange, fill opacity=0.2
[Commit Message Generation, fill=Orange, fill opacity=0.2
NNGen (Liu et al., 2018), ChangeDoc (Huang et al., 2020), fill=Orange, fill opacity=0.2
Loyola et al. (2017), Jiang et al. (2017),
Loyola et al. (2018), PtrGNCMsg (Liu et al., 2019c),
CoDiSum (Xu et al., 2019a), ATOM (Liu et al., 2022c),
CC2Vec (Hoang et al., 2020), CoRec (Wang et al., 2021b), fill=Orange, fill opacity=0.2
CoreGen (Nie et al., 2021), CommitBERT (Jung, 2021),
RACE (Shi et al., 2022), FIRA (Dong et al., 2022),
COME (He et al., 2023a), ExGroFi (Wang et al., 2023b), fill=Orange, fill opacity=0.2
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2
[Text-to-Text, fill=Dandelion, fill opacity=0.4
@@
[Code Evaluation, fill=black, fill opacity=0.7, text=white
[Text-to-Code, fill=yellow, fill opacity=0.5]
[Code-to-Code, fill=cyan, fill opacity=0.4]
[Code-to-Text, fill=Orange, fill opacity=0.2]
[Code-to-Pattern, fill=VioletRed, fill opacity=0.2
[Type
Prediction, fill=VioletRed, fill opacity=0.2
JSNice (Raychev et al., 2015), TypeDevil (Pradel et al., 2015),
Xu et al. (2016), Pigeon (Alon et al., 2018),
Typpete (Hassan et al., 2018), fill=VioletRed, fill opacity=0.2
DeepTyper (Hellendoorn et al., 2018), Jangda & Anand (2019),
NL2Type (Malik et al., 2019), DLTPy (Boone et al., 2019),
TypeWriter (Pradel et al., 2020), OptTyper (Pandi et al., 2020),
Typilus (Allamanis et al., 2020), LambdaNet (Wei et al., 2020),
Type4Py (Mir et al., 2022), HiTyper (Peng et al., 2022),
PYInfer (Cui et al., 2021), fill=VioletRed, fill opacity=0.2
TypeBert (Jesse et al., 2021), TypeWeaver (Yee & Guha, 2023),
TypeT5 (Wei et al., 2023), OpenTau (Cassano et al., 2023c),
TypeGen (Peng et al., 2023c),
Lucchetti & Guha (2024)
@@
[Defect
Detection, fill=VioletRed, fill opacity=0.2
Ray et al. (2016), Bugram (Wang et al., 2016a),
NAR-Miner (Bian et al., 2018), (Pearce et al., 2022), fill=VioletRed, fill opacity=0.2
Wang et al. (2016b), VulDeePecker (Li et al., 2018b),
Lin et al. (2018a),
DeepBugs (Pradel & Sen, 2018), Russell et al. (2018),
SySeVR (Li et al., 2022i),
Devign (Zhou et al., 2019), Li et al. (2019c),
Lin et al. (2021),
VulDeeLocator (Li et al., 2022h), VulDeePecker (Zou et al., 2021),
ReVeal (Chakraborty et al., 2022c), BugLab (Allamanis et al., 2021),
IVDetect (Li et al., 2021a),
ReGVD (Nguyen et al., 2022), fill=VioletRed, fill opacity=0.2
GREAT (Hellendoorn et al., 2020), VulBERTa (Hanif & Maffeis, 2022),
LineVul (Fu & Tantithamthavorn, 2022),
DeepDevVuln (Chan et al., 2023), Yuan et al. (2023a),
(Zhang et al., 2023a), CausalVul (Rahman et al., 2023),
Gao et al. (2023d),
Ullah et al. (2023),
LLM4Vuln (Sun et al., 2024),
MuCoLD (Mao et al., 2024),
Steenhoek et al. (2024)
@@
[Malware
Detection, fill=VioletRed, fill opacity=0.2
@@
Malbert (Xu et al., 2021b),
SHERLOCK (Seneviratne et al., 2022),
MSDT (Tsfaty & Fire, 2022),
MalBERTv2 (Rahali & Akhloufi, 2023),
SocketAI Scanner (Zahan et al., 2024)
@@
[Clone
Detection, fill=VioletRed, fill opacity=0.2
Deckard (Jiang et al., 2007), SourcererCC (Sajnani et al., 2016),
CCAligner (Wang et al., 2018b), LVMapper (Wu et al., 2020),
SAGA (Li et al., 2020a),
NIL (Nakagawa et al., 2021), fill=VioletRed, fill opacity=0.2
White et al. (2016), CDLH (Wei & Li, 2017),
Oreo (Saini et al., 2018), DeepSim (Zhao & Huang, 2018),
ASTNN (Zhang et al., 2019a),
TBCCD (Yu et al., 2019a), Perez & Chiba (2019)
CLCDSA (Nafi et al., 2019),
FA-AST (Wang et al., 2020d),
Fang et al. (2020), Yahya & Kim (2022), fill=VioletRed, fill opacity=0.2
Khajezade et al. (2022),
SSCD (Chochlov et al., 2022),
CCT-LM (Sorokin et al., 2023),
Dou et al. (2023),
Yuan et al. (2023a),
ZC3 (Li et al., 2023d), fill=VioletRed, fill opacity=0.2
[Code
Reasoning, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
[MMLU (Hendrycks et al., 2021b),
CodeQA (Liu & Wan, 2021),
CS1QA (Lee et al., 2022),
C-Eval (Huang et al., 2023b),
CMMLU (Li et al., 2023b),
CodeApex (Fu et al., 2023), fill=VioletRed, fill opacity=0.2
[Code/Author Classification, fill=VioletRed, fill opacity=0.2
SCAP (Frantzeskou et al., 2011), Mahbub et al. (2022), fill=VioletRed, fill opacity=0.2
[Bandara & Wijayarathna (2013),
TBCNN (Mou et al., 2016), inst2vec (Ben-Nun et al., 2018),
DL-CAIS (Abuhamad et al., 2018),
ASTNN (Zhang et al., 2019a), InferCode (Bui et al., 2021a), fill=VioletRed, fill opacity=0.2
Puri et al. (2021), Wang et al. (2022a), fill=VioletRed, fill opacity=0.2
[Machine Code
Detection, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
DetectGPT4Code (Yang et al., 2023e), fill=VioletRed, fill opacity=0.2
[Text-to-Text, fill=Dandelion, fill opacity=0.4
[Document Translation, fill=Dandelion, fill opacity=0.4
[, fill=Dandelion, fill opacity=0.4
[, fill=Dandelion, fill opacity=0.4
[Lu et al. (2021), fill=Dandelion, fill opacity=0.4]
@@
Analysis, fill=Dandelion, fill opacity=0.4
@@
Logram (Dai et al., 2022), SPINE (Wang et al., 2022e), fill=Dandelion, fill opacity=0.4
@@
UniParser (Liu et al., 2022e), fill=Dandelion, fill opacity=0.4
@@
LogPrompt (Liu et al., 2023l), LogGPT (Qi et al., 2023),
Mudgal & Wouhaybi (2023), LLMParser (Jiang et al., 2023d), fill=Dandelion, fill opacity=0.4
2.1.3 Code-to-Text
Code-to-text tasks take code as input, and output text. Related works are listed in Figure 6.
- Code summarization, also referred to as docstring generation, aims to generate a natural language description for a given piece of code (often a function or method). This is the opposite of code synthesis, and SMT/NMT techniques have been likewise applied. Zhang et al. (2022) provide a survey on this topic.
@@ 2.1.3 Code-to-Text
- Identifier prediction is the task of predicting identifier names in the code. As these names are deemed to contain important semantic information, this task has been utilized for code summarization (Allamanis et al., 2016b), as well as pretraining code models (Wang et al., 2021e; Niu et al., 2022). A special case of identifier prediction is method name prediction.

### v6

@@ 2.1.2 Development
Maniple (Parasaram et al., 2024), fill=Orange, fill opacity=0.2
[Development:
Code Suggestion, fill=Orange, fill opacity=0.2
[Type
Prediction, fill=Orange, fill opacity=0.2
JSNice (Raychev et al., 2015), TypeDevil (Pradel et al., 2015),
Xu et al. (2016), Pigeon (Alon et al., 2018),
Typpete (Hassan et al., 2018), fill=Orange, fill opacity=0.2
DeepTyper (Hellendoorn et al., 2018), Jangda & Anand (2019),
NL2Type (Malik et al., 2019), DLTPy (Boone et al., 2019),
TypeWriter (Pradel et al., 2020), OptTyper (Pandi et al., 2020),
Typilus (Allamanis et al., 2020), LambdaNet (Wei et al., 2020),
Type4Py (Mir et al., 2022), HiTyper (Peng et al., 2022),
PYInfer (Cui et al., 2021), fill=Orange, fill opacity=0.2
TypeBert (Jesse et al., 2021), TypeWeaver (Yee & Guha, 2023),
TypeT5 (Wei et al., 2023), OpenTau (Cassano et al., 2023c),
TypeGen (Peng et al., 2023c),
Lucchetti & Guha (2024)
, fill=Orange, fill opacity=0.2
[Identifier Prediction, fill=Orange, fill opacity=0.2
JSNice (Raychev et al., 2015), Pigeon (Alon et al., 2018),
HeMa (Jiang et al., 2019), fill=Orange, fill opacity=0.2
Allamanis et al. (2015), Allamanis et al. (2016b),
GGNN (Allamanis et al., 2018), Code2Vec (Alon et al., 2019b),
Code2Seq (Alon et al., 2019a), Fernandes et al. (2019),
HIER (Xu et al., 2019b), Liu et al. (2019a),
MNire (Nguyen et al., 2020), DeepName (Li et al., 2021b),
DMACOS (Xie et al., 2021b), NamPat (Yang et al., 2022), fill=Orange, fill opacity=0.2
GTNM (Liu et al., 2022b), GTrans (Kuang et al., 2022),
Mario (Wang et al., 2023e), fill=Orange, fill opacity=0.2
[Cloze Test, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
CodeBERT (Feng et al., 2020),
Lu et al. (2021),
Puri et al. (2021),
JavaBERT (de Sousa & Hasselbring, 2021), fill=Orange, fill opacity=0.2
[Development:
Code Explanation, fill=Orange, fill opacity=0.2
[Comment
Generation, fill=Orange, fill opacity=0.2
AutoComment (Wong et al., 2013),
CloCom (Wong et al., 2015)
, fill=Orange, fill opacity=0.2
DeepCom (Hu et al., 2018a),
APIContext2Com (Shahbazi & Fard, 2023)
, fill=Orange, fill opacity=0.2
LAMNER (Sharma et al., 2022),
GTrans (Kuang et al., 2022),
DOME (Mu et al., 2023a),
Geng et al. (2024),
MESIA (Pan et al., 2024)
, fill=Orange, fill opacity=0.2
[Code Summarization, fill=Orange, fill opacity=0.2
, fill=Orange, fill opacity=0.2
CODE-NN (Iyer et al., 2016),
TL-CodeSum (Hu et al., 2018b), Code2Seq (Alon et al., 2019a),
Fernandes et al. (2019), Wan et al. (2018),
AST-AttendGRU (LeClair et al., 2019), Wei et al. (2019),
Haque et al. (2020), DMACOS (Xie et al., 2021b),
Bansal et al. (2021),
CoCoSUM (Wang et al., 2021e),
MLCS (Zhou et al., 2023c), fill=Orange, fill opacity=0.2
TranS3 (Wang et al., 2020e), Ahmad et al. (2020),
Corder (Bui et al., 2021b), SiT (Wu et al., 2021),
SG-Trans (Gao et al., 2023c),
Codex-D (Chen et al., 2021b),
M2TS (Gao & Lyu, 2022), AST-Trans (Tang et al., 2022),
CoSS (Shi et al., 2023a),
Sun et al. (2023b), Yuan et al. (2023a),
use-seq (Su & McMillan, 2023a),
Su & McMillan (2023b),
CSA-Trans (Oh & Yoo, 2024),
Haldar & Hockenmaier (2024)
, fill=Orange, fill opacity=0.2
[Code Documentation, fill=Orange, fill opacity=0.2
, fill=Orange, fill opacity=0.2
, fill=Orange, fill opacity=0.2
Khan & Uddin (2022),
CodeExp (Cui et al., 2022),
HotGPT (Su et al., 2023b),
Dvivedi et al. (2023)
, fill=Orange, fill opacity=0.2
[Document Translation, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
[, fill=Orange, fill opacity=0.2
[Lu et al. (2021), fill=Orange, fill opacity=0.2]
[Testing &
Analysis, fill=cyan, fill opacity=0.2]
[Deployment & Operations, fill=VioletRed, fill opacity=0.2]
[LLM Evaluation, fill=Dandelion, fill opacity=0.4]
Code Editing
- Code translation aims to translate a piece of code (usually a function or method) into another programming language. The relation between code translation and cross-lingual code search is similar to the one between program synthesis and text-to-code retrieval, and SMT/MNT models have also been widely applied to this task. One of the important applications of code translation is migrating old projects written in obsolete languages. However, we are yet to witness such applications at scale in the LLM era, as the context window of even the most powerful language models are quite limited in the face of such projects. Malyala et al. (2023) provide a short survey on this task from the SE perspective.
Code Suggestion
- Type prediction aims to predict the type of variables in dynamic programming languages such as Python and JavaScript. It has been used as a pretraining objective for code language models (Wang et al., 2022e), where it is often simplified as a binary tagging task to predict which tokens in the code are identifiers (Wang et al., 2021d; f).
- Identifier prediction is the task of predicting identifier names in the code. As these names are deemed to contain important semantic information, this task has been utilized for code summarization (Allamanis et al., 2016b), as well as pretraining code models (Wang et al., 2021f; Niu et al., 2022). A special case of identifier prediction is method name prediction.
Code Explanation
- Code summarization, comment generation, and code documentation all aim to generate explanations for code to facilitate understanding of the code, but with slightly different emphasis. Code summarization generates a natural language summary of the given code, while comment generation generates comments. These two terms are often used interchangeably in the literature, and their intended audience is usually developers. Code documentation, on the other hand, is more structured, includes more detailed information about the code’s usage and functions, and is usually intended for a broader audience. Zhang et al. (2022) provide a survey on code summarization.
2.1.3 Testing and Analysis
The third stage in software development is about testing and analyzing the correctness of programs (as well as other properties, such as security). Related works are listed in Figure 6.
Testing
- Unit test generation aims to generate unit tests for a given program. Prior to the rise of Codex and other code LLMs, almost all works in this area employed non-neural methods (see Figure 6). In the age of LLMs, however, this task is ever more important, as research has shown that the current unit tests for evaluating LLMs’ program synthesis capability may be insufficient (Liu et al., 2023f). Wang et al. (2023b) provide a comprehensive survey on software testing with LLMs.
@@ 2.1.3 Testing and Analysis
where level=2text width=3.5em,font=,
where level=3text width=10.5em,font=,
where level=4text width=5.7em,font=,
where level=5text width=10.8em,font=,
@@
[NLP in SE, fill=black, fill opacity=0.7, text=white
[Requirement, fill=yellow, fill opacity=0.5]
[Development, fill=Orange, fill opacity=0.2]
[Testing, fill=cyan, fill opacity=0.2
@@
DeepREL (Deng et al., 2022c), NNSmith (Liu et al., 2023e),
@@
[Analysis, fill=cyan, fill opacity=0.2
[Defect
Detection, fill=cyan, fill opacity=0.2
Ray et al. (2016), Bugram (Wang et al., 2016a),
NAR-Miner (Bian et al., 2018), (Pearce et al., 2022)
, fill=cyan, fill opacity=0.2, text width=6em
Wang et al. (2016b), VulDeePecker (Li et al., 2018b),
Lin et al. (2018a),
DeepBugs (Pradel & Sen, 2018), Russell et al. (2018),
SySeVR (Li et al., 2022j),
Devign (Zhou et al., 2019), Li et al. (2019c),
Lin et al. (2021),
VulDeeLocator (Li et al., 2022i), VulDeePecker (Zou et al., 2021),
ReVeal (Chakraborty et al., 2022c), BugLab (Allamanis et al., 2021),
IVDetect (Li et al., 2021a),
ReGVD (Nguyen et al., 2022)
, fill=cyan, fill opacity=0.2, text width=10.2em
GREAT (Hellendoorn et al., 2020), VulBERTa (Hanif & Maffeis, 2022),
LineVul (Fu & Tantithamthavorn, 2022),
DeepDevVuln (Chan et al., 2023), Yuan et al. (2023a),
(Zhang et al., 2023a), CausalVul (Rahman et al., 2023),
Gao et al. (2023d),
Ullah et al. (2023),
LLM4Vuln (Sun et al., 2024),
MuCoLD (Mao et al., 2024),
Steenhoek et al. (2024)
, fill=cyan, fill opacity=0.2
[Clone
Detection, fill=cyan, fill opacity=0.2
Deckard (Jiang et al., 2007), SourcererCC (Sajnani et al., 2016),
CCAligner (Wang et al., 2018c), LVMapper (Wu et al., 2020),
SAGA (Li et al., 2020a),
NIL (Nakagawa et al., 2021)
, fill=cyan, fill opacity=0.2, text width=6em
White et al. (2016), CDLH (Wei & Li, 2017),
Oreo (Saini et al., 2018), DeepSim (Zhao & Huang, 2018),
ASTNN (Zhang et al., 2019a),
TBCCD (Yu et al., 2019a), Perez & Chiba (2019)
CLCDSA (Nafi et al., 2019),
FA-AST (Wang et al., 2020d),
Fang et al. (2020), Yahya & Kim (2022)
, fill=cyan, fill opacity=0.2, text width=10.2em
Khajezade et al. (2022),
SSCD (Chochlov et al., 2022),
CCT-LM (Sorokin et al., 2023),
Dou et al. (2023),
Yuan et al. (2023a),
ZC3 (Li et al., 2023d)
, fill=cyan, fill opacity=0.2
[Malware
Detection, fill=cyan, fill opacity=0.2
, fill=cyan, fill opacity=0.2, text width=6em
, fill=cyan, fill opacity=0.2, text width=10.2em
Malbert (Xu et al., 2021b),
SHERLOCK (Seneviratne et al., 2022),
MSDT (Tsfaty & Fire, 2022),
MalBERTv2 (Rahali & Akhloufi, 2023),
SocketAI Scanner (Zahan et al., 2024)
, fill=cyan, fill opacity=0.2
[Code/Author Classification, fill=cyan, fill opacity=0.2
SCAP (Frantzeskou et al., 2011), Mahbub et al. (2022)
, fill=cyan, fill opacity=0.2, text width=6em
[Bandara & Wijayarathna (2013),
TBCNN (Mou et al., 2016), inst2vec (Ben-Nun et al., 2018),
DL-CAIS (Abuhamad et al., 2018),
ASTNN (Zhang et al., 2019a), InferCode (Bui et al., 2021a)
, fill=cyan, fill opacity=0.2, text width=10.2em
Puri et al. (2021), Wang et al. (2022a), fill=cyan, fill opacity=0.2
[Deployment & Operations, fill=VioletRed, fill opacity=0.2]
[LLM Evaluation, fill=Dandelion, fill opacity=0.4]
Analysis
- Malware detection is a similar task to defect detection. However, malware differs from other types of vulnerable code in that the bugs therein are malicious - i.e. they are intentionally injected. We refer to Sahin & Bahtiyar (2020) and Gopinath & Sethuraman (2023) for reviews on non-Transformer based methods for this task.
2.1.4 Deployment and Operations
After software passes the testing stage, it is deployed into production, and often continuously maintained or updated. For this stage, we summarize the specific tasks and their related works in Figure 7.
@@ 2.1.4 Deployment and Operations
[NLP in SE, fill=black, fill opacity=0.7, text=white
[Requirement, fill=yellow, fill opacity=0.5]
[Development, fill=Orange, fill opacity=0.2]
[Testing &
Analysis, fill=cyan, fill opacity=0.2]
[Deployment, fill=VioletRed, fill opacity=0.2
[Compiler
Optimization, fill=VioletRed, fill opacity=0.2
[, fill=VioletRed, fill opacity=0.2
DeepTune (Cummins et al., 2017)
@@
Armengol-Estapé & O’Boyle (2021),
Cummins et al. (2023),
Priority Sampling (Grubisic et al., 2024b),
Grubisic et al. (2024a)
@@
[Decompilation, fill=VioletRed, fill opacity=0.2
@@
BTC (Hosseini & Dolan-Gavitt, 2022),
SLaDe (Armengol-Estapé et al., 2023),
DecGPT (Wong et al., 2023),
Nova+ (Jiang et al., 2023c),
LLM4Decompile (Tan et al., 2024)
@@
[Deobfuscation, fill=VioletRed, fill opacity=0.2
DeGuard (Bichsel et al., 2016), Autonym (Vasilescu et al., 2017),
Debin (He et al., 2018), JSNeat (Tran et al., 2019), fill=VioletRed, fill opacity=0.2
DIRE (Lacomis et al., 2019), Artuso et al. (2021), fill=VioletRed, fill opacity=0.2
VarBERT (Banerjee et al., 2021), DIRECT (Nitin et al., 2021),
LmPa (Xu et al., 2023), fill=VioletRed, fill opacity=0.2
@@
Analysis, fill=VioletRed, fill opacity=0.2
@@
Logram (Dai et al., 2022), SPINE (Wang et al., 2022f)
, fill=VioletRed, fill opacity=0.2
@@
UniParser (Liu et al., 2022e)
, fill=VioletRed, fill opacity=0.2
@@
LogPrompt (Liu et al., 2023k), LogGPT (Qi et al., 2023),
Mudgal & Wouhaybi (2023), LILAC (Jiang et al., 2023d),
LLMParser (Ma et al., 2024)
, fill=VioletRed, fill opacity=0.2
[Operations, fill=VioletRed, fill opacity=0.2
[Code
Review, fill=VioletRed, fill opacity=0.2
Review Bot (Balachandran, 2013), Tricorder (Sadowski et al., 2015),
ClusterChanges (Barnett et al., 2015), CommentFinder (Hong et al., 2022), fill=VioletRed, fill opacity=0.2
DCR (Gupta & Sundaresan, 2018), DeepReview (Li et al., 2019a),
DACE (Shi et al., 2019), CORE (Siow et al., 2020), fill=VioletRed, fill opacity=0.2
Tufano et al. (2021b), Hellendoorn et al. (2021),
Tufano et al. (2022b), CodeReviewer (Li et al., 2022k),
AUGER (Li et al., 2022c), crBERT (Yin et al., 2023),
LLaMA-Reviewer (Lu et al., 2023a), fill=VioletRed, fill opacity=0.2
[Workflow
Automation, fill=VioletRed, fill opacity=0.2
, fill=VioletRed, fill opacity=0.2
, fill=VioletRed, fill opacity=0.2
Mehta et al. (2023),
Zhang et al. (2024c)
, fill=VioletRed, fill opacity=0.2
[Commit Message Generation, fill=VioletRed, fill opacity=0.2
NNGen (Liu et al., 2018), ChangeDoc (Huang et al., 2020)
, fill=VioletRed, fill opacity=0.2
Loyola et al. (2017), Jiang et al. (2017),
Loyola et al. (2018), PtrGNCMsg (Liu et al., 2019c),
CoDiSum (Xu et al., 2019a), ATOM (Liu et al., 2022c),
CC2Vec (Hoang et al., 2020), CoRec (Wang et al., 2021c)
, fill=VioletRed, fill opacity=0.2
CoreGen (Nie et al., 2021), CommitBERT (Jung, 2021),
RACE (Shi et al., 2022), FIRA (Dong et al., 2022),
COME (He et al., 2023a), ExGroFi (Wang et al., 2023c)
, fill=VioletRed, fill opacity=0.2
[LLM Evaluation, fill=Dandelion, fill opacity=0.4]
Deployment
Compiler optimization is the task of optimizing compiled code to increase its efficiency. Many early works applied reinforcement learning and other machine learning technologies to this task by optimizing task scheduling or pass ordering (Wang & O’Boyle, 2018; Leather & Cummins, 2020), and it is only recently that researchers started to view it as a sequence-to-sequence generation task with the help of powerful LLMs (Cummins et al., 2023).
- Decompilation is the reverse process of compilation, and an important topic in reverse engineering. In this task a model takes a compiled program - represented in assembly code or binary machine code - and aims to output the original source program in high-level languages such as C.
- Obfuscation refers to the process of renaming identifiers (e.g. variables, methods, and classes), for example to generic names like var_1, var_2 or x, y. It is an important technique in virus detection, intellectual property protection, and code size reduction (Collberg & Thomborson, 2002; Murad et al., 2010; Vasilescu et al., 2017). Deobfuscation refers to the reverse process, where meaningful identifier names are recovered from obfuscated programs. Obfuscation can be easily achieved statically, but deobfuscation has been a subject of more interest in recent years. It plays a significant role in decompiling, and has also been adopted as a pretraining objective for code language models (Lachaux et al., 2021; Ding et al., 2022a; Liu et al., 2022d).
Operations